OpenStudy (anonymous):

A park maintenance person stands 19 m from a circular monument. If you assume her lines of sight form tangents to the monument and make an angle of 34°, what is the measure of the arc of the monument that her lines of sight intersect? 146, 56, 124, or 112?

OpenStudy (anonymous):

 If the segments are tangent, then there are two arcs on the circle. The large one away from the person and the smaller one the person is looking at. Let x = the small arc Then the bigger arc is 360 - x since there are 360 degrees in a circle. Recall the theorem: an angle outside of a circle is equal to half the difference of the two arcs (something like that). Using this, we can set up the equation: 34 = (1/2)((360 - x) - x) because the angle on the angle formed by her lines of sight 34 = (1/2)(360 - 2x) 34 = (180 - x) x = 146 There's your answer! Dunno why you were given 19 meters... Hopefully I didn't mess up!